ShaquilleOatmeal1:

Ok so i got a question on my worksheet and it says 2^3 * 2^5 and i got 2^8 is that right?

AZ:

Correct! Because \(\Large a^b \times a^c = a^{b+c}\)

ShaquilleOatmeal1:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @AZ Correct! Because \(\Large a^b \times a^c = a^{b+c}\) \(\color{#0cbb34}{\text{End of Quote}}\) Ok thank you :D i just wanted to make sure

AZ:

Of course! You're welcome :)

DontQuestionMe:

Lmao we are both working on exponential functions how weird the answer is correct just like how az said

ShaquilleOatmeal1:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @DontQuestionMe Lmao we are both working on exponential functions how weird the answer is correct just like how az said \(\color{#0cbb34}{\text{End of Quote}}\) LMAOOOOOOOOOOOOOOOOOOOO thx bella XD